A circle has a diameter of 10 centimeters and a central angle EOG that measures 210 degrees. What is the area of the shaded sector? Use 3.14 for pi and round your answer to the nearest tenth.
A. 2.6 cm^2
B. 5.2 cm^2
C. 45.8 cm^2
D. 51.5 cm^2

Answer by lwsshak3(11628)

A circle has a diameter of 10 centimeters and a central angle EOG that measures 210 degrees. What is the area of the shaded sector? Use 3.14 for pi and round your answer to the nearest tenth.
area of sector=1/2*r^2*central angle(radians)=1/2*5^2*7pi/6=45.8 cm^2
note: 210 ˚=(210/180)pi radians=7pi/6 radians